Skip navigation

Gravity Probe B

Testing Einstein's Universe


The Payload Test II Phase A testing has gone very well during the past week. The four gyroscope readout systems have been tested and have demonstrated functionality. This test shows that the replaced gyroscope #4 now meets readout system functionality. The tests performed include basic tune up of the SQUIDs, DC coupling of the gyroscope to the SQUID as demonstrated by gyroscope vibration induced SQUID noise, successful feedback to the input circuit of the SQUID, and measurement of high frequency SQUID noise measured to 100 kHz.

The telescope test team verified end-to-end functionality of the flight telescope system. This test consisted of injecting light into the telescope aperture and collecting data at the output of the room temperature electronics. The test confirmed that the telescope and the four optical channels (sensed by 2 cryogenic detectors each) function as expected. This test shows that both the primary and backup telescope systems are fully functional.

The GMA valves, plumbing, and bottles have been transferred to the flight plate. The last step in the assembly of GMA, the installation of instrumentation and wiring onto the GMA is in-progress. Following assembly, the GMA will begin acceptance testing.

Photo: An O-ring on one of the many metering valves of the Gas Management Assembly. Click here or on the small photo to read a detailed description of the gyro #4 replacement earlier this year.

The GP-B team successfully performed a data flow test between the ITF Pod B and a Wallops Flight Facility (WFF) programmable telemetry processor (PTP). The test verified the command and telemetry interfaces between the pod ground software and the NASA Ground Network PTP.

Raw 2.56 Mbps telemetry data was decoded, derandomized, and Reed-Solomon error corrected by the WFF PTP. The PTP then stripped out the realtime telemetry and sent it to Pod B via a TCP/IP socket. The data also included playback data which was recorded at the PTP as required. Commands were also sent from Pod B via another socket connection and their receipt was verified by the WFF enigineer.

The forward and aft ECU units passed functional test and were random vibed normal to the baseplate. Post vibe test is underway. Two temperature measurements are out of spec; indications point to the test set. Verification of the out of tolerance readings will commence upon completion of the full post vibe test.

The SRE team is awaiting release by incoming inspection of the aft unit B converter boards, which have been conformally coated. Four of the five DAS boards in the forward SREs are through inspection. The fifth board failed incoming inspection and is being returned for rework; this board should be available next week. It is anticipated that aft unit B (unit A has already been through ATP) and the two forward units will be assembled by the end of next week.

Photo: LM engineers Peter Carley and Bill Given in the ITF in LM B/205.

<<Previous Highlight   Next Highlight>>

Index of Highlights